Three numbers are in the ratio 2:3:4. What are the numbers if their sum is 72?

Answer by checkley79(3341)

2N+3N+4N=9N
9N=72
N=72/9
N=8 ANS. FOR THE MULTIPLIER.
2*8=16 ANS. FOR THE SMALLER NUMBER.
3*8=24 ANS. FOR THE SECOND NUMBER.
4*8=32 ANS. FOR THE LARGER NUMBER.
PROOF:
16+24+32=72
72=72
2/16=3/24=4/32
1/8=1/8=1/8